OpenStudy (anonymous):

Factor completely and then place the factors in the proper location on the grid.25 a2 - 30 a + 9

OpenStudy (anonymous):

the question is 25a^2-30a+9 right?

OpenStudy (anonymous):

if that is the case, your factoring should look like this: (5a-3)(5a-3) because -3*-3 =9(right hand term) and -3(5a)+-3(5a) =-15a-15a =-30a (middle term) and 5a*5a=25a^2

OpenStudy (anonymous):

not sure about the grid though. do they give you a grid?